# ]organization=LASIGE, Faculdade de Ciências, Universidade de Lisboa , country=Portugal

### 26 Jan 2024 v1

@@ 3.3 Chatbots for Cybersecurity
The reviewed papers in this section primarily focused on evaluating and testing the commercial ChatGPT model, with limited attention given to assessing open-source chatbot models in the context of cybersecurity applications. Moreover, to the best of our knowledge, there is a notable absence of comprehensive comparative studies explicitly dedicated to open-source chatbots within the specialized field of cybersecurity. This gap in existing research underscores the need to comprehensively examine open-source chatbot models, their effectiveness, and their potential contributions to enhancing cyber threat awareness and detection. Furthermore, researchers have yet to evaluate chatbots for two specific downstream tasks in the cybersecurity domain. However, for the NER task in the public domain, the ChatGPT model has been evaluated [30]. Our study aims to fill this gap by carefully comparing open-source chatbots in the context of a CTI application, considering two NLP tasks.
@@ 4 LLM-based Chatbots
LLM-based chatbots simulate human-like conversations with users through text or speech interaction. Chatbots can offer users more intelligent and contextually relevant responses to queries by utilizing LLMs’ language generation and comprehension capabilities. LLM-powered chatbots have grown in popularity in various fields such as customer service, virtual assistants, and information retrieval. This section reviews commercial and open-source chatbot models, providing a comparative analysis of eight chatbot models, namely, LLaMA, GPT4all, Dolly 2.0, Stanford Alpaca, Alpaca-LoRA, Vicunna, Falcon, and ChatGPT. We aim to assess their effectiveness in detecting cybersecurity-related texts and identifying relevant entities. By examining their capabilities in these specific domains, we aimed to gain insights into the strengths and weaknesses of each model.
@@ 5 Chatbots Evaluation
We divided this section into three topics relevant to the evaluation: dataset, methodology, and evaluation criteria. First, we introduce the Twitter dataset, which serves as the foundational source for generating subsequent prompts. The methodology subsection outlines the techniques and strategies used to assess and compare chatbot performances. Finally, in the evaluation criteria subsection, we explore the metrics and standards by which the chatbots are assessed.
@@ 5.1 Dataset
We leveraged a comprehensive Twitter dataset made available by Alves et al. [4] in their work to retrieve the Indicators of Compromise (IoCs) from Twitter OSINT. This dataset contains a combined total of 31281 tweets collected during two distinct periods: one from November 21, 2016, to March 27, 2017, and the other from June 1, 2018, to September 1, 2018. After collection, the tweets were filtered using specific keywords and manually labeled as positive or negative, considering their relevance to the cybersecurity of an IT infrastructure, thus creating labeled datasets suitable for supervised learning. The dataset consists of the tweet release time, tweet content, cleaned tweet, label, and entity columns.
We extracted information from the cleaned tweet column to create a customized dataset for questions and prompts. This process yielded a new dataset of 31281 constructed questions. This set of questions provides a foundation for binary classification and NER experiments, enabling the evaluation of chatbot performance in these tasks.
@@ 5.2 Methodology
Our primary objective is to test the capabilities of chatbots to address two critical NLP tasks: binary classification and NER in Twitter OSINT. We interacted with the chatbots using Python scripts to iterate through the dataset, presenting each question to the chatbot’s prompt, and recording the obtained responses.
@@ 5.2.1 Binary Classification
We designed prompts to determine whether tweets are related to the cybersecurity field, aiming to generate a response including ‘yes’ or ‘no’. We focused on constraining the chatbot to provide concise yes or no answers without additional explanations. This restriction simplifies the process of extracting binary labels (0 or 1) from the responses. For instance, in a scenario where the “clean tweet" column within our dataset contained the text “cyber infosec kenno media SQL injection", to compose the specific prompts we desired, we added the phrase “Is the sentence" at the beginning of each clean tweet. Then, we appended the phrase “related to cybersecurity? Please respond with yes or no." at the end. Consequently, the final prompt becomes: Is the sentence ‘cyber infosec kenno media SQL injection’ related to cybersecurity? Please respond with either yes or no.
This process was applied to all the tweets in the “Clean Tweet" column to create a question dataset. We administered three distinct tests, each consisting of the 31281 questions asked to chatbots.
@@ 5.2.2 Named Entity Recognition
The initial approach, which is called Entity-Specific Prompting (ESP), was chosen based on a careful evaluation of the various experimental results. It was evident from these experiments that attempting to extract all required entities from a single tweet using a single question yielded significant misrecognition, rendering the approach impractical. Consequently, we decided on a more precise strategy involving the creation of specific prompts for each entity within each tweet. After sending several prompt requests for each entity to the chatbots, we found that the organization names and product version entities were the most extractable by chatbots. The revised approach focuses on two specific entities: organization names (B-ORG) and product versions (B-VER). We selected these entities for extraction by interacting with ChatGPT-3.5-turbo, ChatGPT-4, GPT4all, and Dolly models. Since only 11074 out of the total 31281 questions in the dataset were tagged with NER labels, we limited our scope to the initial 2550. In our experiment, we employed various prompts to choose the proper one; ultimately, we selected the following prompts to identify the organization name and product version:
@@
The second approach, which is called Guide-Line Prompting (GLP), exclusively employed for ChatGPT-4, involves a comprehensive specification of all entities within the ‘GUIDELINES_PROMPT’ section. This approach attempts to extract seven entities in a single prompt and considers only the ChatGPT model because the GUIDELINES_PROMPT feature is absent in open-source models. In this guideline, we include two examples of tweets from 11074 NER-tagged tweets in the dataset, each annotated with their respective entities. In addition, we have included the output format as a means to direct ChatGPT’s response for subsequent processing. We then sent a dedicated prompt for each of the 11074 tweets and systematically covered all the extracted entities. The prompt guideline is shown in Fig. 1.
Fig. 1: Second NER approach: employing a ChatGPT-4 template for entity identification on tweets
@@ 5.3 Evaluation Criteria
In this section, we examine the evaluation criteria used to assess chatbot models’ performance in addressing binary classification and NER tasks. Our evaluation methodology focuses on performance and quality.
@@ 5.3.1 Performance
To assess the binary classification tasks, we used the F1-score, which considers the balance between the True Positive Rate (TPR) and True Negative Rate (TNR). Similarly, in the case of the NER tasks, we computed the F1-score by considering Precision and Recall.
@@ 5.3.2 Quality
By analyzing these distinct response modes from the output files, we gain valuable insights into the performance and capabilities of the evaluated chatbot models to effectively detect and address cybersecurity-related questions. The experimental results section provides a detailed explanation of the experimental findings, focusing on the evaluation criteria of quality and accuracy.
@@ 6.1 Adherence to Proper Grammar and Spelling
The fundamental principle behind creating prompts is ensuring clear and precise instructions. It is crucial to express expectations and provide explicit instructions. These approaches enhance the possibility of directing the model to the desired output while decreasing the chances of receiving irrelevant or incorrect responses. It is important to note that writing a clear prompt should not be confused with writing a short prompt. Longer prompts have the potential to provide increased clarity and context to the model, thereby resulting in more comprehensive and relevant responses. Thus, finding the right balance between being brief and providing chatbot models with sufficient information can make the prompt more effective[46]. A useful tactic to help us write clear and specific prompts is to use separators or markers to clearly show the different parts of the input.
We devoted considerable attention to formulating suitable prompts or questions to ensure accurate and relevant responses from the chatbot models during our experiment. The objective was to design prompts that effectively captured the essence of the cybersecurity-related information present in the tweets. By formulating questions such as: Is the sentence ‘vuln oracle java se cve-2016-5582 remote security vulnerability’ related to cybersecurity? Just answer yes or no. We aimed to elicit yes or no responses from chatbots. Our technique involves generating prompts, querying the chatbot models, analyzing the output, and iterating this process multiple times until we achieve the desired prompt formulation. Using spaces and adequately employing the apostrophe (’) played a significant role in clarifying the prompt.
We iteratively evaluated the chatbot models’ responses based on different prompts, seeking those that yielded the most precise responses, particularly those with yes and no. This allowed us to assess the chatbot models’ capability to detect cybersecurity concepts and their ability to generate meaningful and contextually appropriate responses. Once we have formulated the desired prompt, we automate the process of sending the entire set of questions to the chatbots.
@@ 6.2 Text Length Control
The parameter that defines the length of the answer or the number of tokens in local chatbots is represented as ‘N_predict’. As the number of questions in the dataset increases, the time required for these models to provide answers also increases. To optimize the execution time, we advise setting a parameter that controls the response length to the smallest suitable value based on a specific task. We consistently set the value of N_predict to 15 across all compared models. This parameter plays a critical role in significantly reducing the execution time for answering questions across all chatbot models.
In the context of ChatGPT-3.5-turbo and ChatGPT-4, utilizing the max_tokens parameter constrains the extent of the response generated by the model. This is achieved by establishing a predetermined upper limit for the number of tokens that can be words or characters within the generated output. Utilization of this parameter enabled us to control the length of the responses. Using more extended responses in ChatGPT-4 can increase token consumption, potentially increasing the usage costs. The parameter value 70 was assigned to the ChatGPT-3.5-turbo and ChatGPT-4 models in various experimental tests. It is worth noting that the chosen token length of 15 for the open-source chatbots applies exclusively to the generated answers. By contrast, the selected token length of 70 for ChatGPT encompasses questions and answers.
@@ 7 Experimental Results
It is essential to highlight that we ultimately elaborate on the common and default parameters shared by all open-source chatbot models. In most of these models, the maximum size of the context window and its default value are set to 512. An exception is the Dally model, which has a maximum context window size of 1024. However, the commercial ChatGPT model exhibits variability across its final versions, each with a distinct context window size. For instance, ChatGPT-4 is available in two versions with window sizes of 8k and 32k, whereas ChatGPT-3.5-turbo is available in 4k and 16k versions. In our experiments, we utilized a server equipped with multiple GPU units, including an NVIDIA A30 GPU (memory capacity: 24,576 MiB) and an NVIDIA RTX A6000 GPU (memory capacity: 49,140 MiB), with a total system memory of 263,951,816 kB.
@@ 7.1.1 Performance
In terms of accuracy, we explain the results of the tests conducted on chatbot models. Table 1 provides an overview of the tests conducted for each chatbot and its version. Moreover, we recorded the execution time of the models’ responses for all the questions. It includes details regarding the number of parameters on which the model was trained, achieved F1-scores, and execution time for each model. By analyzing the F1-score values in Test 1, we can assess the effectiveness of the models in accurately responding to the given questions. Based on the results presented in Table 1, it is evident that the GPT4all achieved the highest accuracy among the open-source models, as indicated by its F1-score of 0.90. The Dolly model has an accuracy of 0.86, Falcon of 0.85, Alpaca-LoRA of 0.84, and the Stanford Alpaca model has a score of 0.64. Although GPT4all achieved higher accuracy among open-source chatbots, it is noteworthy that the commercial ChatGPT models (GPT-4 and GPT-3.5-turbo) achieved an F1-score of 0.94. ChatGPT-3.5-turbo with a 16k window context size achieves the same F1-score as ChatGPT-4 with an 8k context window size. These results highlight the better accuracy of the GPT4all and ChatGPT models, emphasizing their effectiveness for this particular task. The Fig. 3 depicts a summary of the data presented in Table 1, specifically highlighting the F1-Scores of chatbots with the highest number of parameters in the context of Test 1.
@@
Based on the shuffled dataset test, the GPT4all model achieved an F1-score of 0.89%, whereas the Dolly model attained an F1-score of 0.86%.
@@
This indicates a slight decrease in accuracy of approximately 1% for the GPT4all model compared with the first test, which can be considered insignificant. It is essential to mention that shuffling the prompt does not affect the accuracy of the Falcon, Stanford Alpaca, and Alpaca-LoRA models, as the F1-score is equal to that of the first test.
@@
Upon applying the isolated prompt test to the dataset, the F1-score for the GPT4all model attained a value of 0.88. This test results in a 2% decrease compared with that of Test 1. Conducting the test for the Stanford Alpaca model did not result in a significant reduction in accuracy. When testing the Alpaca-LoRA model, we observed a two percent reduction in the F1-score of 0.82. The result of this test on the model Falcon 40B is two percent less than the first test, and it is equal to 0.83. We ignored this test on Dolly because it was time-consuming and beyond the available time and resources. It is worth noting that running this test increases the execution time of the models. This is because of the need to launch the model for each question. Tests 2 and 3 were excluded from the ChatGPT model because of the high cost of utilizing their API. Nevertheless, the results obtained by the other models indicate that we should not expect a significant decrease in accuracy.
@@
Chatbots are based on a varying number of parameters. This study involved models ranging from 7B to 65B in parameter counts. Based on our experiment, the number of parameters significantly influences the effectiveness of the models in answering questions. Generally, models with 7B parameters demonstrated lower quality and could not provide definitive yes or no answers. Remarkably, GPT4all, with 7B parameters, faced limitations in providing yes or no responses to every question. In response to each question, the output consisted solely of explanations, leading to a significant expenditure on human effort to discern whether the response was yes or no, thus preventing the automated processing of the answers. Consequently, we were unable to calculate the F1-score for the 7B parameter model. In addition, two other models, Dolly and Stanford Alpaca, both of which have precisely 7B parameters, have lower F1-scores than the models with 13B parameters, as shown in Table 1.
Model Test Number
Trained Parameters
F1-score Execution Time
ChatGPT-3.5-turbo (16k context) [6] Test 1 175B 0.9431 11h 23m
ChatGPT-4 (8k context) [6] Test 1 1.7T 0.9410 11h 50m
GPT4all [7] Test 1 13B 0.9049 132h 05m
GPT4all Test 2 13B 0.8927 132h 02m
GPT4all Test 3 13B 0.8844 136h 05m
Dolly 2.0 [8] Test 1 7B 0.8470 10h 38m
Dolly 2.0 Test 1 12B 0.8612 10h 16m
Dolly 2.0 Test 2 12B 0.8631 10h 00m
Dolly 2.0 Test 3 12B - LET
Falcon [11] Test 1 7B 0.8304 16h 02m
Falcon Test 1 40B 0.8511 54h 03m
Falcon Test 2 40B 0.8502 54h 55m
Falcon Test 3 40B 0.8330 71h 10m
Alpaca-LoRA [10] Test 1 65B 0.8477 10h 12m
Alpaca-LoRA Test 2 65B 0.8451 10h 44m
Alpaca-LoRA Test 3 65B 0.8241 11h 20m
Stanford Alpaca [9] Test 1 7B 0.3112 13h 03m
Stanford Alpaca Test 1 13B 0.4209 13h 21m
Stanford Alpaca Test 1 30B 0.6415 15h 48m
Stanford Alpaca Test 2 30B 0.6401 15h 04m
Stanford Alpaca Test 3 30B 0.6395 16h 18m
Vicuna [12] Test 1 13B 0.3611 11h 23m
Dionisio et al. [5] Test 1 - 0.9470 00h 43m
* LET: Long Execution Time
Fig. 3: F1-score of using chatbots for cybersecurity text classification - Test 1

### 13 Mar 2024 v2

@@ 2.7 The Research Gap
The reviewed papers in this section primarily focused on evaluating and testing the commercial ChatGPT model, with limited attention given to assessing open-source chatbot models in the context of cybersecurity applications. Moreover, to the best of our knowledge, there is a notable absence of comprehensive comparative studies explicitly dedicated to open-source chatbots within the specialized field of OSINT-based CTI. This gap in existing research underscores the need to examine open-source chatbot models, their effectiveness, and their potential contributions to enhancing cyber threat awareness and detection. Although there are specialized state-of-the-art models, including DL models that excel at CTI binary classification and NER tasks [11, 41, 10], no comparative study has been conducted to determine whether LLM-based chatbots can compete with their performance. If they can provide competitive results, CTI tools could be changed to integrate LLM chatbots into the OSINT processing pipeline, decreasing the tools’ complexity and maintenance costs. By employing general-purpose LLM chatbots, tasks related to CTI data collection, curation, labelling, and model training and updating would no longer be as necessary as they are for specialized models.
Our study aims to fill this gap by carefully comparing open-source and publicly available paid chatbots in the context of an OSINT-based CTI application, considering two downstream NLP tasks: binary classification and NER.
@@ 3 LLM-based chatbots
LLM-based chatbots simulate human-like conversations with users through text or speech interaction. They offer users more intelligent and contextually relevant responses to queries by utilizing LLMs’ language comprehension and generation capabilities. This section reviews the eight state-of-the-art commercial and open-source chatbot models available at the time of writing, namely, LLaMA, GPT4all, Dolly 2.0, Stanford Alpaca, Alpaca-LoRA, Vicunna, Falcon, and ChatGPT. We used these models in our experimental evaluation (Section 6) to assess their effectiveness in detecting cybersecurity-related texts and identifying relevant entities.
@@ 4 Chatbots evaluation
We divided this section into three topics relevant to the evaluation: dataset, methodology, and evaluation criteria. First, we introduce the Twitter dataset, which is the foundational source for generating subsequent prompts. The methodology subsection outlines the techniques and strategies used to assess and compare chatbot performances. Finally, in the evaluation criteria subsection, we explore the metrics and standards by which the chatbots are assessed.
@@ 4.1 Dataset
We leveraged a comprehensive Twitter dataset made available by Alves et al. [12] in their work to retrieve the Indicators of Compromise (IoCs) from Twitter OSINT. This dataset contains a combined total of 31281 tweets collected during two distinct periods: one from November 21, 2016, to March 27, 2017, and the other from June 1, 2018, to September 1, 2018. After collection, the tweets were filtered using specific keywords and manually labelled as positive or negative, considering their relevance to the cybersecurity of an IT infrastructure, thus creating labelled datasets suitable for supervised learning. Later, Dionísio et al. labeled the dataset [11] for NER and republished it.
The dataset consists of 31281 tweet records, including the timestamp of the tweet, specific keywords found in the tweet, the original tweet, a pre-processed tweet cleaned from some special characters, a binary label marking the tweet as relevant for cybersecurity or not, and a string identifying the named entities in the pre-processed tweet. Table B1 in Appendix B shows representative examples of dataset records.
In this work, we used the pre-processed tweets, the cybersecurity relevance, and the sequences of named entity tags, to create a customized dataset for the binary classification and NER experiments through which the chatbots will be evaluated.
@@ 4.2 Methodology
As discussed in previous work [10, 11, 41], the extraction of actionable CTI from OSINT entails three foundational tasks: text classification, extraction of pertinent information from textual sources, and the synthesis of the information gathered into concise summaries. Considering the role of summarization to aggregate correlated threat indicators prior to dissemination, this study directs its attention towards evaluating and comparing chatbot performance within the two first tasks. These NLP tasks, being more amenable to LLM methodologies [31], are deemed as primary focal points for research within the scope of this study.
From a practical point of view, we interacted with the chatbots using Python language scripts to iterate through the dataset entries. For each entry, different prompts were used to address the two tasks of interest: deciding on the relevance of a tweet for cybersecurity using a binary classification formulation and extracting relevant information from the tweets using NER.
@@ 4.2.1 Binary classification
We designed prompts to determine whether tweets are related to the cybersecurity field, aiming to generate a response including ‘yes’ or ‘no’. We focused on constraining the chatbot to provide concise yes or no answers without additional explanations. This restriction simplifies the process of extracting binary labels (0 or 1) from the responses. For instance, in a scenario where the “pre-processed tweet" column within our dataset contained the text “cyber infosec kenno media SQL injection", to compose the specific prompts we desired, we added the phrase “Is the sentence" at the beginning of each pre-processed tweet. Then, we appended the phrase “related to cybersecurity? Please respond with yes or no." at the end. Consequently, the final prompt becomes: Is the sentence ‘threatmeter dos microsoft internet explorer 9 mshtml cdisp node::insert sibling node use-after-free ms13-0’ related to cybersecurity? Just answer yes or no.
This process was applied to all the tweets in the “pre-processed tweet" column to create a question dataset. We administered three distinct tests, each consisting of the 31281 questions asked to chatbots.
@@ 4.2.2 Named entity recognition
The initial approach, which is called Entity-Specific Prompting (ESP), was chosen based on a careful evaluation of the various experimental results. It was evident from these experiments that attempting to extract all required entities from a single tweet using a single question yielded significant misrecognition, rendering the approach impractical. Consequently, we decided on a more precise strategy involving the creation of specific prompts for each entity within each tweet. After sending several prompt requests for each entity to the chatbots, we found that the organization names and product version entities were the most extractable by chatbots. The revised approach focuses on two specific entities: organization names (B-ORG) and product versions (B-VER). We selected these entities for extraction by interacting with ChatGPT-3.5-turbo, ChatGPT-4, GPT4all, and Dolly models. Since only 11074 out of 31281 questions in the dataset had been tagged with NER labels, we limited our analysis for the NER task to these 11074 tweets. In our experiment, we employed various prompts to choose the proper one; ultimately, we selected the following prompts to identify the organization name and product version:
@@
The second approach, which is called Guide-Line Prompting (GLP), exclusively employed for ChatGPT-4, involves a comprehensive specification of all entities within the ‘GUIDELINES_PROMPT’ section. This approach attempts to extract seven entities in a single prompt and considers only the ChatGPT model because the GUIDELINES_PROMPT feature is absent in open-source models. In this guideline, we include two examples of tweets from 11074 NER-tagged tweets in the dataset, each annotated with their respective entities. In addition, we have included the output format as a means to direct ChatGPT’s response for subsequent processing. We then sent a dedicated prompt for each of the 11074 tweets and systematically covered all the extracted entities. The prompt guideline is given in Appendix A.
@@ 4.3 Evaluation criteria
This section describes our evaluation criteria to assess chatbot models’ performance in addressing binary classification and NER tasks. Our evaluation methodology focuses on performance and quality.
@@ 4.3.1 Performance
To assess the selected chatbots’ performance, we used the $\text{F}_{1}$ score, which is a metric that computes the harmonic mean of Precision and Recall to measure the binary classification performance of a model. It strikes a balance between the proportion of true positive results in all positive predictions and the proportion of true positive results in all actual positives. The $\text{F}_{1}$ score is given by
$\text{F}_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (1)
where Precision is the ratio of correctly predicted positive observations to the total predicted positives, given by
$\text{Precision}=\frac{TP}{TP+FP},$
and Recall measures the ratio of correctly predicted positive observations to all observations in the positive class:
$\text{Recall}=\frac{TP}{TP+FN}.$
@@ 4.3.2 Quality
By analyzing these distinct response modes from the output files, we gain valuable insights into the performance and capabilities of the evaluated chatbot models to effectively detect and address cybersecurity-related questions. The experimental results section provides a detailed explanation of the experimental findings, focusing on the quality and accuracy evaluation criteria.
@@ 5.1 Adherence to proper grammar and spelling
The fundamental principle behind creating prompts is ensuring clear and precise instructions. Prompt engineering [64] is crucial for optimizing the performance of chatbot models by enhancing the clarity and specificity of the given instructions. Thorough prompt design and testing improve the ability of the model to comprehend requests, making it a more effective tool for generating desired outcomes. These approaches enhance the possibility of directing the model to the desired output while decreasing the chances of receiving irrelevant or incorrect responses.
We devoted considerable attention to formulating suitable prompts to maximize accurate and relevant responses from the chatbots during our experiments. The objective was to design prompts that effectively allow models to capture the essence of the CTI task in the information present in tweets.
Following the best practices for prompt engineering [58], we progressively refined our prompts using two approaches. For NER, we considered first a prompt guideline template approach but with poor results (prompt template is shown in Appendix A). We used a zero-shot methodology for binary classification, considering that LLMs are excellent zero-shot reasoners [65]. Since this approach produced very good results, for the NER tests, we transformed the prompt guideline into a sequence of zero-shot questions, one for each entity we aim to extract.
To leverage the zero-shot capability of LLMs, we explored the assumption that chatbots perfectly represent the cybersecurity concept. Following this assumption, we designed prompts by starting the question with information on the cybersecurity issue and finishing it with precise instructions on the expected outcome.
The process of selecting the final prompt was iterative and relied on a few trial and error cycles, using the principles described before. Using a sample of dataset entries, we progressively queried the chatbots and evaluated the answers until consistent answers were achieved concerning the instructions given. This refinement cycle resulted in the determination of the final prompt.
By formulating questions such as Is the sentence ‘vuln oracle java se cve-2016-5582 remote security vulnerability’ related to cybersecurity? Just answer yes or no., we aimed to elicit yes or no responses from the chatbots. Using spaces and adequately employing the apostrophe (’) played a significant role in clarifying the prompt.
This methodology allowed us to assess the chatbot’s capabilities to detect cybersecurity concepts and their ability to generate meaningful and contextually appropriate responses. Having formulated the desired prompt, we automated the process of sending the entire set of questions to chatbots.
@@ 5.2 Text length control
The parameter that defines the length of the answer or the number of tokens in local chatbots is represented as N_predict. It plays a critical role in significantly reducing the execution time for answering questions across all open-source chatbot models. To optimize the execution time, we advise setting a parameter that controls the response length to the smallest suitable value based on a specific task. After some initial experimentation, we consistently set the value of N_predict to 15 across all open-source models.
In the context of ChatGPT-3.5-turbo and ChatGPT-4, the max_tokens parameter constrains the length of the responses generated by the model. This is achieved by establishing a predetermined upper limit for the number of tokens that can be words or characters within the generated output. Using more extended responses in ChatGPT-4 can increase token consumption, potentially increasing the usage costs. After some initial experimentation with prompts with extreme lengths, the value 70 was assigned to max_tokens in the experimental tests. It is worth noting that the chosen token length of 15 for the open-source chatbots applies exclusively to the generated answers. By contrast, the selected token length of 70 for ChatGPT models encompasses questions and answers.
@@ 6 Results and Discussion
It is essential to highlight that we ultimately elaborate on the common and default parameters shared by all open-source chatbot models. In most of these models, the maximum size of the context window [^3] Upper limit for the range of tokens considered to provide an answer. and its default value are set to 512. An exception is the Dolly model, which has a maximum context window size of 1024. However, the commercial ChatGPT model exhibits variability across its final versions, each with a distinct context window size. For instance, ChatGPT-4 is available in two versions with window sizes of 8k and 32k, whereas ChatGPT-3.5-turbo is available in 4k and 16k versions. In our experiments, we utilized a server equipped with multiple GPU units, including an NVIDIA A30 GPU (memory capacity: 24,576 MiB) and an NVIDIA RTX A6000 GPU (memory capacity: 49,140 MiB), with 264 GB of RAM.
@@ 6.1 Evaluation of binary classification
Model Test Number
Parameters
Precision Recall $\text{F}_{1}$ score Execution Time
ChatGPT-3.5-turbo (16k context) [13] Test 1 175B 0.9570 0.9280 0.9431 11h 23m
ChatGPT-3.5-turbo (16k context) [13] Test 2 175B 0.9700 0.9200 0.9489 11h 23m
ChatGPT-3.5-turbo (16k context) [13] Test 3 175B - - UECH -
ChatGPT-4 (8k context) [13] Test 1 1.7T 0.9580 0.9240 0.9410 11h 50m
ChatGPT-4 (8k context) [13] Test 2 1.7T 0.9590 0.9230 0.9403 11h 43m
ChatGPT-4 (8k context) [13] Test 3 1.7T - - UECH -
GPT4all [14] Test 1 13B 0.9490 0.8630 0.9049 132h 05m
GPT4all Test 2 13B 0.9490 0.8410 0.8927 132h 02m
GPT4all Test 3 13B 0.9470 0.8280 0.8844 136h 05m
Dolly 2.0 [15] Test 1 7B 0.8890 0.8000 0.8470 10h 38m
Dolly 2.0 Test 1 12B 0.9470 0.7900 0.86120 10h 16m
Dolly 2.0 Test 2 12B 0.9480 0.7910 0.8631 10h 00m
Dolly 2.0 Test 3 12B - - - LET
Falcon [18] Test 1 7B 0.8120 0.8500 0.8304 16h 02m
Falcon Test 1 40B 0.8980 0.8200 0.8511 54h 03m
Falcon Test 2 40B 0.8990 0.8080 0.8502 54h 55m
Falcon Test 3 40B 0.8990 0.7880 0.8330 71h 10m
Alpaca-LoRA [17] Test 1 65B 0.8980 0.7940 0.8477 10h 12m
Alpaca-LoRA Test 2 65B 0.8990 0.8000 0.8451 10h 44m
Alpaca-LoRA Test 3 65B 0.8980 0.7610 0.8241 11h 20m
Stanford Alpaca [16] Test 1 7B 0.2260 0.5000 0.3112 13h 03m
Stanford Alpaca Test 1 13B 0.3240 0.6000 0.4209 13h 21m
Stanford Alpaca Test 1 30B 0.6980 0.6050 0.6415 15h 48m
Stanford Alpaca Test 2 30B 0.6990 0.5920 0.6401 15h 04m
Stanford Alpaca Test 3 30B 0.6990 0.5810 0.6395 16h 18m
Vicuna [19] Test 1 13B 0.4390 0.3100 0.3611 11h 23m
Dionisio et al.  [41] Test 1 - 0.9570 0.9363 0.9470 00h 43m
* LET: Long Execution Time                   * UECH : Uncertainty of Erasing Conversation History
@@ 6.1.1 Performance
In terms of accuracy, we explain the results of the tests conducted on chatbot models. Table 1 provides an overview of the tests conducted for each chatbot and its version. Moreover, we recorded the execution time of the models’ responses for all the questions. It includes details regarding the number of parameters on which the model was trained, achieved $\text{F}_{1}$ score, precision, recall, and execution time for each model. The confusion matrices corresponding to each model are provided in Appendix C. By analyzing the $\text{F}_{1}$ score values in Test 1, we can assess the effectiveness of the models in accurately responding to the given questions. Based on the results presented in Table 1, it is evident that the GPT4all achieved the highest accuracy among the open-source models, as indicated by its $\text{F}_{1}$ score of 0.90. The Dolly model has an accuracy of 0.86, Falcon of 0.85, Alpaca-LoRA of 0.84, and the Stanford Alpaca model has a score of 0.64. Although GPT4all achieved higher accuracy among open-source chatbots, it is noteworthy that the commercial ChatGPT models (GPT-4 and GPT-3.5-turbo) achieved an $\text{F}_{1}$ score of 0.94. ChatGPT-3.5-turbo with a 16k window context size achieves the same $\text{F}_{1}$ score as ChatGPT-4 with an 8k context window size. These results highlight the better accuracy of the GPT4all and ChatGPT models, emphasizing their effectiveness for this particular task.
@@
Based on the shuffled dataset test, the GPT4all model achieved an $\text{F}_{1}$ score of 0.89%, whereas the Dolly model attained an $\text{F}_{1}$ score of 0.86%.
@@
This indicates a slight decrease in accuracy of approximately 1% for the GPT4all model compared with the first test, which can be considered insignificant. It is essential to mention that shuffling the prompt does not affect the accuracy of the ChatGPT, Falcon, Stanford Alpaca, and Alpaca-LoRA models, as the $\text{F}_{1}$ score is equal to that of the first test.
@@
Upon applying the isolated prompt test to the dataset, the $\text{F}_{1}$ score for the GPT4all model attained a value of 0.88. This test results in a 2% decrease compared with that of Test 1. Conducting the test for the Stanford Alpaca model did not result in a significant reduction in accuracy. When testing the Alpaca-LoRA model, we observed a two percent reduction in the $\text{F}_{1}$ score of 0.82. The result of this test on the model Falcon 40B is two percent less than the first test, and it is equal to 0.83. We ignored this test on Dolly because it was time-consuming and beyond the available time and resources. For ChatGPT, we used the available API to send requests. In this case, Test 3 is not feasible because isolated prompt functionality is not available, i.e., the conversation history of the model cannot be completely reset, and we have no means to reinitialize the model for each prompt. It is worth noting that running this test increases the execution time because of the need to launch the model for each question. Nevertheless, the results obtained by the other models indicate that we should not expect a significant decrease in accuracy.
@@
Chatbots are based on a varying number of parameters. This study involved models ranging from 7B to 65B in parameter counts. Based on our experiment, the number of parameters significantly influences the effectiveness of the models in answering questions. Generally, models possessing 7B parameters exhibited a comparatively lower performance [66] and, based on our experiments, failed to provide binary (’yes’ or ’no’) answers. Remarkably, GPT4all, with 7B parameters, faced limitations in providing yes or no responses to every question. In response to each question, the output consisted solely of explanations, leading to a significant expenditure on human effort to discern whether the response was yes or no, thus preventing the automated processing of the answers. Consequently, we were unable to calculate the $\text{F}_{1}$ score for the 7B parameter model. In addition, two other models, Dolly and Stanford Alpaca, both of which have precisely 7B parameters, have lower $\text{F}_{1}$ score than the models with 13B parameters, as shown in Table 1.